OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 72 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 34, how many women auditioned? 22 44 50 60

OpenStudy (anonymous):

m + w = 72 --> m = 72 - w 2m - 3w = 34 Now just sub 72 - w in for m in the second equation and you will find the number of women that auditioned. Can you do that ?

OpenStudy (anonymous):

Once you sub it in, solve for w :)

OpenStudy (anonymous):

is it 44 w

OpenStudy (anonymous):

let me check....hold on

OpenStudy (anonymous):

or is it 50 w and 22 m

OpenStudy (anonymous):

2m - 3w = 34 2(72 - w) - 3w = 34 144 - 2w - 3w = 34 -5w = 34 - 144 -5w = - 110 w = -110/-5 w = 22 m + w = 72 m + 22 = 72 m = 72 - 22 m = 50 check... 2m - 3w = 34 2(50) - 3(22) = 34 100 - 66 = 34 100 = 100 (correct) It is 22 women and 50 men Any questions ?

OpenStudy (anonymous):

thank u very much

OpenStudy (anonymous):

your very welcome :)